### v1

@@ 4 Application
We consider the shallow water equations on the one dimensional spatial domain $S=[-L_{S}/2,L_{S}/2]$ with periodic boundary conditions. The system of equations governs the evolution of water height $h(t,\cdot)$ and momentum $q(t,\cdot)=h(t,\cdot)v(t,\cdot)$, with $v(t,\cdot)$ denoting the velocity field and $t\geq 0$ time. The scalar fields $h$ and $q$ are assumed to take values in a sufficiently regular subset of $L^{2}(S,\nu;\mathbb{R})$. The governing equations read
(10) $\partial_{t}h+\partial_{s}q$ $=0$
$\partial_{t}q+\partial_{s}(q^{2}/h+1/2Fr^{-2}h^{2})$ $=0$
subject to an appropriate initial condition. In the above $\partial_{t}$ and $\partial_{s}$ denote differentiation with respect to time and space respectively; $Fr$ denotes the flow’s Froude number, which is a dimensionless ratio of inertial over gravitational forces. Using the state variable $x=(h,q)\in X\subset L^{2}(S,\nu;\mathbb{R}^{2})$ and function $f(x)=(q,q^{2}/h+1/2Fr^{-2}h^{2})$ we rewrite Eq. 10 in the compact form
(11) $\partial_{t}x+\partial_{s}f(x)=0.$
@@
for all $t\geq 0$. If one assumes that the height and velocity fields are differentiable for all $t\geq 0$, thereby excluding discontinuous solutions, then total energy is conserved along solution trajectories, namely $dE(t)/dt=0$. In this case the state space can be decomposed into the dynamics invariant disjoint subsets
$\mathcal{E}_{c}:=\{x\in X:E(t)=c\text{ for all $t\geq 0$}\}$
@@
In the formation of discontinuities, the internal energy of the fluid is increased across a discontinuity, accompanied by an increase in entropy. Since the above definition of total energy does not include internal energy, total energy is decreased across a discontinuity [42]. In the present work we consider only differentiable solutions.
@@ 4.1 Discretization
We employ a finite volume spatial discretization of Eq. 11 using a uniform fine grid of $M_{f}\in\mathbb{N}$ cells with length $\Delta s=L/M_{f}$ each, centered at points $s_{m}\in S$, $m\in\{0,\ldots,M_{f}-1\}$. We denote $S_{f}=\{s_{m}\}_{m=0}^{M_{f}-1}\subset M$ and $C_{m}=[s_{m}-\Delta s/2,s_{m}+\Delta s/2]\subset S$, $m\in\{0,\ldots,M_{f}-1\}$. Each cell $C_{m}$ can also be represented by its center point $s_{m}\in S_{f}$. We define the cell averaged state variable
@@
reusing the same state symbol $x$. For any $t\geq 0$ function $x(t,\cdot)$ can be represented by vector $\underline{x}\in\mathbb{R}^{2M_{f}}$ with $x_{m}=x(t,s_{m})$, $s_{m}\in C_{m}$.
We derive the semidiscretized system of equations
(12) $\dot{x}_{m}=\frac{1}{\Delta s}(F_{m}-F_{m+1})$
using the local Lax-Friedrich (LLF) numerical flux
(13) $F_{m}=\frac{1}{2}[f(x_{m-1})+f(x_{m})]-\frac{\lambda_{m}}{2}(x_{m}-x_{m-1})$
@@
(14) $\lambda_{m}=\max\,\bigl{(}\lvert q_{m-1}/h_{m-1}\rvert+Fr^{-1}\sqrt{h_{m-1}},\,\lvert q_{m}/h_{m}\rvert+Fr^{-1}\sqrt{h_{m}}\bigr{)},$
where subscripts denote components of the associated vectors. In the above $F_{m}$ denotes the numerical flux at the left face of cell $C_{m}$ [25]. Based on Eq. 13 and Eq. 14 we define the formal map $\mathcal{F}_{LLF}$
$F_{m}=\mathcal{F}_{LLF}(x_{m-1},x_{m}).$
We now form a coarse grid and associated coarse variables defined as local spatial averages of the fine variables $x_{m}$. In particular we form $M_{c}\in\mathbb{N}$ coarse grid cells by merging every $M_{s}\in\mathbb{N}$ fine cells; in this way, $M_{c}=M_{f}/M_{s}$ and each coarse cell has length $\Delta\hat{s}=M_{s}\Delta s$. We define the resulting coarse grid $S_{c}=\{\hat{s}_{m}\}_{m=0}^{M_{c}-1}\subset S$ and cells $\hat{C}_{m}=[\hat{s}_{m}-\Delta\hat{s}/2,\hat{s}_{m}+\Delta\hat{s}/2]\subset S$, $m\in\{0,\ldots,M_{c}-1\}$. The coarse grid state variable $\hat{x}$ is formed using spatial averages over each coarse cell
$\hat{x}(t,s)=\frac{1}{M_{s}}\sum_{j=M_{s}m}^{M_{s}(m+1)-1}x_{j},\quad s\in\hat{C}_{m}.$
For the evolution of the coarse state variable $\hat{x}$ we derive the system of equations
(15) $\dot{\hat{x}}_{m}=\frac{1}{\Delta\hat{s}}[(\hat{F}_{m}+G_{m})-(\hat{F}_{m+1}+G_{m+1})]$
@@
By definition, the coarse grid fluxes Eq. 16 depend only on the coarse variables. On the contrary, the subgrid fluxes Eq. 17 depend on both the coarse and the original fine variables; in particular, they depend on the values of $x_{m}$ in the fine cells that border the interfaces of each coarse cell. The subgrid fluxes act as antidiffusive corrections to the coarse grid fluxes, increasing the accuracy of the discretization compared to one performed entirely on the coarse grid [39].
@@ 4.2 Closure model
The state space of the true spatiotemporal dynamics (11) is an infinite dimensional Hilbert space $X\subset L^{2}(S,\nu;\mathbb{R}^{2})$. To be able to simulate the “true” dynamics numerically we approximate the full state space $X\subset L^{2}(S,\nu_{f};\mathbb{R}^{2})$ by a Hilbert subspace of the space of $L^{2}$ functions on the mesh produced by our fine discretization. Next we define the resolved state space $\mathcal{X}=L^{2}(S,\nu_{c};\mathbb{R}^{2})\cap X$ and decompose the full state space as $X=\mathcal{X}\oplus\mathcal{Y}$ with unresolved state space $\mathcal{Y}=\mathcal{X}^{\perp}$. In this way the resolved state space can be considered as the space of piecewise constant functions, each taking a constant value on each cell formed by the coarse discretization. We set flux space $\mathcal{Z}=L^{2}(S,\nu_{c};\mathbb{R}^{2})$ and true flux map $Z:\mathcal{Y}\rightarrow\mathcal{Z}$ with action given by (17). Now the full dynamics $\Phi^{n\Delta t}:X\rightarrow X$ is given by the discrete time version of (12) and the resolved dynamics $\Phi_{\mathcal{X}}^{n\Delta\hat{t}}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{X}$ by the discrete time version of (15), with $n\in\mathbb{N}$ and timesteps $\Delta t>0$ and $\Delta\hat{t}=M_{s}\Delta t$.
@@
Using the full dynamics we generate $N$ samples of the resolved state $\{\hat{x}_{n}\}_{n=0}^{N-1}\subset\mathcal{X}$ and fluxes $\{z_{n}\}_{n=0}^{N-1}\subset\mathcal{Z}$ which may consist of one or more dynamical trajectories. Each sample $\hat{x}_{n}$ or $z_{n}$ is represented by a vector in $\mathbb{R}^{2M_{c}}$ holding the values corresponding to the height $h$ and momentum $q$ variables in each coarse grid cell. We define the product state space $\Omega:=X\times S$ and associated Hilbert space $H_{\Omega,N}:=L^{2}(\Omega,\sigma_{N};\mathbb{R})$ with $\sigma_{N}=\mu_{N}\times\nu_{M}$ and $\nu_{M}=\nu_{c}$. For notational consistency with the previous sections, in the following we denote by $M=M_{c}$ the number of coarse grid cells and by $S_{M}=S_{c}\subset S$ the set of grid points representing the coarse grid cells.
@@
To form our closure model we define the surrogate unresolved state space $\tilde{\mathcal{Y}}=L^{2}(S,\nu_{M};B(H_{\Omega}))$ with associated surrogate flux map $\tilde{Z}:\tilde{\mathcal{Y}}\rightarrow\mathcal{Z}$ and surrogate resolved dynamics $\tilde{\Phi}_{\mathcal{X}}^{n\Delta\hat{t}}:\mathcal{X}\times\tilde{\mathcal{Y}}\rightarrow\mathcal{X}$. To define the action of the surrogate flux map $\tilde{Z}$ we introduce two quantum observables $A^{(j)}\in B(H_{\Omega,N})$, $j\in\{0,1\}$. Observable $A^{(0)}$ will be used to produce the flux values for the height $h$ variables; observable $A^{(1)}$ the flux values for the momentum $q$ variables. With these choices, the fluxes for the $h$ variables are computed as $\tilde{z}^{(0)}(s_{m})=\operatorname{tr}(\rho(s_{m})A^{(0)})$ and those for the $q$ variables as $\tilde{z}^{(1)}(s_{m})=\operatorname{tr}(\rho(s_{m})A^{(1)})$, $s_{m}\in S_{M}$.
@@
In Appendix A we explain the sequence of steps used to build the QMCL scheme and apply it to predict the dynamics based on a given initial condition.
@@ 4.3 Numerical results
(18) $h_{0}(s)$ $=1+0.3(1-\delta/2)\sin\bigl{(}\frac{2\pi}{L_{S}}3.5(1-\delta/2)s+\pi/6\bigr{)}$
$v_{0}(s)$ $=1+0.2(1-\delta)\sin\bigl{(}\frac{2\pi}{L_{S}}3(1-\delta)s\bigr{)}$
@@
with $s\in S$, $L_{S}=100$, parameter $0\leq\delta\leq 1$, $h_{0}$ the initial condition for height and $q_{0}=h_{0}v_{0}$ the initial condition for momentum. Our training data consists of three trajectories generated using the values $\delta\in\{0,0.5,1\}$. For each trajectory we integrate the full dynamics (12) in time for 12,200 timesteps and then sample the next 3,260 timesteps. Each sample contains the values of functions $h$ and $q$ on the fine grid.
@@
To derive the resolved state samples that will be used as our training data, we coarsen the model and above sampled data. By averaging the sampled data over every $M_{s}=20$ fine cells we arrive at a coarse grid of $M=96$ cells of length $\Delta\hat{s}=20\Delta s$ each. Given that each cell is now $20$ times longer we also increase the sampling timestep by the same amount to set $\Delta\hat{t}=20\Delta t$. From the original 3,260 samples per trajectory generated on the fine grid we keep every 20th sample, leading to 163 samples per trajectory. For each one of those samples we spatially average the solution over every $20$ cells to derive our resolved state samples, namely the averaged values of $h$ and $q$ over the coarse grid. We then derive the subgrid flux training samples using (17).
Using $Q=64$ delays, these samples yield 100 training samples per trajectory in delay embedded form, adding up to $N=300$ samples in total. Each resolved state sample consists of $1922M=$ 192 values, 96 for each of the height and momentum variables. With this training dataset, the dimension of the product Hilbert space $H_{\Omega,N}$ is equal to $NM=28800$. Using the approach described in Section 3.6 for dealing with multitrajectory training data, we build an orthonormal basis using a total of $L=6144$ eigenfunctions and define $H_{\Omega,N}^{L}$ as their linear span. Each eigenfunction is represented by a vector in $\mathbb{R}^{NM}$. Finally, we set $J=5$ to define data map $W_{J}$ that will be used for the bayesian conditioning of quantum states.
@@
To form the orthonormal basis we must compute the eigenvectors of an $NM\times NM$ symmetric kernel matrix $\bm{K}_{N}$ representing the data driven integral operator (8). To reduce the computational cost of this operation we perform a partial Cholesky factorization of $\bm{K}_{N}$ to approximate it by lower rank matrix $\tilde{\bm{K}}_{N}=\bm{F}\bm{F}^{T}$ with $\bm{F}\in\mathbb{R}^{NM\times r}$ and $r<NM$. We perform the factorization using rank parameter $r=L=6144$ and the randomized algorithm developed in [9]. We then use the approximate kernel matrix to perform its bistochastic normalization and solve the eigenvalue problem with reduced cost [40].
@@
To test the predictive performance of the QMCL scheme we use two initial conditions obtained from (18) for $\delta\in\{0.25,0.75\}$, which are trajectories not included in the training dataset. After integrating the true dynamics (12) for 12,200 timesteps we spatially average the found solutions to generate the initial conditions that will be used to initialize the resolved state in QMCL. Using those initial conditions we integrate in time the coarse dynamics (15) using timestep $\Delta\hat{t}$ and the subgrid fluxes predicted by QMCL. In what follows we present results obtained after integrating for a total of 120 timesteps while performing bayesian conditioning of the quantum states every 10 timesteps.
@@
The results comparing the evolution under the true and surrogate dynamics are grouped into four figures. Figures 1 and 3 include the results for the trajectory generated by the initial condition obtained from (18) for $\delta=0.25$. Figures 2 and 4 showcase the results obtained for $\delta=0.75$.
@@
Looking at the heatmaps for the height and momentum fields in Figure 1 for the trajectory with $\delta=0.25$ and Figure 2 for the one with $\delta=0.75$, we see that QMCL clearly captures the main features of the dynamics for the tested time horizon. For both trajectories, QMCL correctly predicts the main traveling wave features of the dynamics, including the spatial variations of the waves as they interact with one another. Moreover, the magnitude of the predicted fields is consistent with the true dynamics in most regions of the spatial domain. In addition to the results shown in these figures, we have also verified that QMCL can reproduce the dynamical trajectories included in its training dataset with good accuracy.
@@
In the present closure model, the subgrid fluxes play the role of antidiffusive corrections to the fluxes generated entirely on the coarse grid [39]. Their values are proportional to the spatial derivative of the associated field, taking large values when the height or momentum field is changing rapidly in space, and approximately zero value when the fields are constant in space. As a result, the contribution of the subgrid fluxes is critical in regions of large spatial variation of the height and momentum fields, counterbalancing the increased diffusion of the solution introduced by our coarsening of the original dynamics by spatial averaging. This implies that when the exact value of the subgrid fluxes is not predicted correctly by QMCL, the surrogate evolution will be more diffusive than the true evolution.
@@
This effect can be seen in the heatmaps for the two trajectories. By comparing the true and predicted fields of height and momentum, we see that the surrogate dynamics leads to stronger diffusion of the solution in areas where there is strong spatial variation. The increased diffusion is seen in the plots as bleeding of color between different space-time regions, with the boundaries between those regions being less sharp than in the true dynamics. Similarly, the same effect can be seen by comparing the subgrid flux fields for height and momentum. In the true dynamics, the peaks of the fluxes are sharp and clearly separated from the valleys, with clearly visible waves traveling in opposite directions. In the predicted dynamics, although the waves are still predicted correctly, their peaks are not as sharp, with their values diffusing more strongly to the surrounding region.
@@
The flux plots also clearly demonstrate the effect of bayesian conditioning, which for these results is performed every $10$ timesteps. One can see the discontinuities that occur every $10$ timesteps for the predicted fluxes over large regions of the spatial domain. These discontinuities represent corrections introduced to the predicted flux values every time conditioning of the quantum states is performed.

### v2

@@ 5 Application
We consider the shallow water equations on the one dimensional spatial domain $S=[-L_{S}/2,L_{S}/2]$ with periodic boundary conditions. The system of equations governs the evolution of water height $h(t,\cdot)$ and momentum $q(t,\cdot)=h(t,\cdot)v(t,\cdot)$ with $v(t,\cdot)$ denoting the velocity field and $t\geq 0$ time. The scalar fields $h(t,\cdot)$ and $q(t,\cdot)$ are assumed to take values in a sufficiently regular subset of $L^{2}(S,\nu;\mathbb{R})$ for all $t\geq 0$. The governing equations read
$\partial_{t}h+\partial_{s}q$ =0=0= 0 (10)
$\partial_{t}q+\partial_{s}(q^{2}/h+1/2Fr^{-2}h^{2})$ =0=0= 0
subject to an appropriate initial condition. In the above $\partial_{t}$ and $\partial_{s}$ denote differentiation with respect to time and space respectively; $Fr$ denotes the flow’s Froude number, which is a dimensionless ratio of inertial over gravitational forces. Using the state variable $x=(h,q)\in X\subset L^{2}(S,\nu;\mathbb{R}^{2})$ and function $f(x)=(q,q^{2}/h+1/2Fr^{-2}h^{2})$ we rewrite (10) in the compact form
$\partial_{t}x+\partial_{s}f(x)=0.$ (11)
@@
for all $t\geq 0$. If one assumes that the height and velocity fields are differentiable for all $t\geq 0$, thereby excluding discontinuous solutions, then the total energy is conserved along solution trajectories, namely $dE(t)/dt=0$. In this case the state space can be decomposed into the dynamics invariant disjoint subsets
$\mathcal{E}_{c}=\{x\in X\colon E(t)=c\text{ for all $t\geq 0$}\}$
@@
In the formation of discontinuities, the internal energy of the fluid is increased across a discontinuity, accompanied by an increase in entropy. Since the above definition of total energy does not include internal energy, the total energy is decreased across a discontinuity [69]. In the present work we consider only differentiable solutions.
@@ 5.1 Discretization
We employ a finite volume spatial discretization of (11) using a uniform fine grid of $M_{f}\in\mathbb{N}$ cells with length $\Delta s=L/M_{f}$ each, centered at points $s_{m}\in S$, $m\in\{0,\ldots,M_{f}-1\}$. We denote $S_{f}=\{s_{m}\}_{m=0}^{M_{f}-1}\subset M$ and $C_{m}=[s_{m}-\Delta s/2,s_{m}+\Delta s/2]\subset S$, $m\in\{0,\ldots,M_{f}-1\}$ where each cell $C_{m}$ can also be represented by its center point $s_{m}\in S_{f}$. We define the cell averaged state variable
@@
reusing the same state symbol $x$. For any $t\geq 0$ function $x(t,\cdot)$ can be represented by the vector $\underline{x}\in\mathbb{R}^{2M_{f}}$ with $x_{m}=x(t,s_{m})$, $s_{m}\in S_{f}$.
We derive the semidiscretized system of equations governing the evolution of the state variables
$\dot{x}_{m}=\frac{1}{\Delta s}(F_{m}-F_{m+1})$ (12)
using the local Lax-Friedrichs (LLF) numerical flux
$F_{m}=\frac{1}{2}[f(x_{m-1})+f(x_{m})]-\frac{\lambda_{m}}{2}(x_{m}-x_{m-1})$ (13)
@@
$\lambda_{m}=\max\,\bigl{(}\lvert q_{m-1}/h_{m-1}\rvert+Fr^{-1}\sqrt{h_{m-1}},\,\lvert q_{m}/h_{m}\rvert+Fr^{-1}\sqrt{h_{m}}\bigr{)}.$ (14)
In the above, subscripts denote components of the associated vectors while $F_{m}$ denotes the numerical flux at the left face of cell $C_{m}$ [79]. Based on (13) and (14) we introduce the notation
$F_{m}=\mathcal{F}_{LLF}(x_{m-1},x_{m})$
which will be used later on to define the coarse grid fluxes.
We now form a coarse grid and associated coarse variables defined as local spatial averages of the fine variables $x_{m}$. In particular we form $M_{c}\in\mathbb{N}$ coarse grid cells by merging every $M_{s}\in\mathbb{N}$ fine cells; in this way, $M_{c}=M_{f}/M_{s}$ and each coarse cell has length $\Delta\hat{s}=M_{s}\Delta s$. We define the resulting coarse grid $S_{c}=\{\hat{s}_{m}\}_{m=0}^{M_{c}-1}\subset S$ with cells $\hat{C}_{m}=[\hat{s}_{m}-\Delta\hat{s}/2,\hat{s}_{m}+\Delta\hat{s}/2]\subset S$, $m\in\{0,\ldots,M_{c}-1\}$. The coarse grid state variable $\hat{x}$ is formed using spatial averages over each coarse cell
$\hat{x}(t,s_{m})=\frac{1}{M_{s}}\sum_{j=M_{s}m}^{M_{s}(m+1)-1}x_{j},\quad s_{m}\in S_{c}$
and is represented by the vector $\underline{\hat{x}}\in\mathbb{R}^{2M_{c}}$ with $\hat{x}_{m}=\hat{x}(t,s_{m})$, $s_{m}\in S_{c}$. For the evolution of the coarse state variable $\hat{x}$ we derive the system of equations
$\dot{\hat{x}}_{m}=\frac{1}{\Delta\hat{s}}[(\hat{F}_{m}+G_{m})-(\hat{F}_{m+1}+G_{m+1})]$ (15)
@@
By definition, the coarse grid fluxes (16) depend only on the coarse variables. On the contrary, the subgrid fluxes (17) depend on both the coarse and the original fine variables; in particular, they depend on the values of $x_{m}$ in the fine cells that border the interfaces of each coarse cell. The subgrid fluxes act as antidiffusive corrections to the coarse grid fluxes, increasing the accuracy of the discretization compared to one performed entirely on the coarse grid [80].
@@ 5.2 Closure model
The state space of the true spatiotemporal dynamics (11) is an infinite dimensional Hilbert space $X\subset L^{2}(S,\nu;\mathbb{R}^{2})$. To be able to simulate the “true” dynamics numerically we approximate the state space by $X\subset L^{2}(S,\nu_{f};\mathbb{R}^{2})$ where $X$ is now considered as a Hilbert subspace of the space of $L^{2}$ functions on the mesh produced by our fine discretization. Next we define the resolved state space $\mathcal{X}=L^{2}(S,\nu_{c};\mathbb{R}^{2})\cap X$ and decompose the full state space as $X=\mathcal{X}\oplus\mathcal{Y}$ with unresolved state space $\mathcal{Y}=\mathcal{X}^{\perp}$. In this way the resolved state space can be considered as a space of piecewise constant functions, each taking a constant value on each cell formed by the coarse discretization. The full dynamics $\Phi^{n\Delta t}\colon X\to X$ is now given by the discrete time version of (12) and the resolved dynamics $\Phi_{\mathcal{X}}^{n\Delta\hat{t}}\colon X\to\mathcal{X}$ by the discrete time version of (15) with $n\in\mathbb{N}$ and timesteps $\Delta t>0$ and $\Delta\hat{t}=M_{s}\Delta t$.
@@
We define the product state space $\Omega=X\times S$ and associated Hilbert space of observables $H_{\Omega,N}=L^{2}(\Omega,\sigma_{N};\mathbb{C})$ with total number of samples $N\in\mathbb{N}$ and measures $\sigma_{N}=\mu_{N}\times\nu_{M}$ and $\nu_{M}=\nu_{c}$. For notational consistency with the previous sections, in the following we denote by $M=M_{c}$ the number of coarse grid cells and by $S_{M}=S_{c}\subset S$ the set of grid points representing the coarse grid cells. We introduce two flux functions $z_{j}\in H_{\Omega,N}$, $j\in\{0,1\}$ producing the fluxes for the height and momentum variables respectively, denoting their collection by $z=\{z_{j}\}_{j=0}^{1}$. For every state $x\in X$ the fluxes $z_{j}(x)$ are given by (17).
Using the full dynamics we generate $N$ samples of the resolved state $\{\hat{x}_{n}\}_{n=0}^{N-1}$ and fluxes $\{z(x_{n},\cdot)\}_{n=0}^{N-1}$ which will be used as our training data and may consist of one or more dynamical trajectories. Each sample $\hat{x}_{n}$ or $z(x_{n},\cdot)$ is represented by a vector in $\mathbb{R}^{2M}$ holding the values corresponding to the height $h$ and momentum $q$ variables in each coarse grid cell.
@@
To form our closure model we define the surrogate unresolved state space $\tilde{\mathcal{Y}}=L^{2}(S,\nu_{M};B(H_{\Omega}))$ and two quantum observables $A_{j}\in B(H_{\Omega,N})$, $j\in\{0,1\}$. Observable $A_{0}$ will be used to produce the surrogate fluxes for the height $h$ variables and is formed as multiplication by the true flux function $z_{0}$; analogously, observable $A_{1}$ will produce the surrogate fluxes for the momentum $q$ variables and is defined as multiplication by $z_{1}$. With these definitions in place, the surrogate fluxes are computed using the two surrogate flux functions $\tilde{z}_{j}\colon\tilde{\mathcal{Y}}\to H_{S}$ with action $\tilde{z}_{0}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A_{0})$ and $\tilde{z}_{1}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A_{1})$ where $\rho\in\tilde{\mathcal{Y}}$ and $s_{m}\in S_{M}$.
@@
In A we explain the sequence of steps used to build the data driven QMCl scheme and apply it to predict the dynamics based on a given initial condition.
@@ 5.3 Numerical results
[…]
with $s\in S$, $L_{S}=50$, parameter $0\leq\delta\leq 1$, $h_{0}$ the initial condition for height and $q_{0}=h_{0}v_{0}$ the initial condition for momentum. The solutions generated by this family of initial conditions feature a number of wave fronts traveling in different directions and interacting with one another. Figures 2 and 3 present space-time heatmaps of the solutions that will be used to test the predictive performance of our closure model.
Figure 1: Space-time heatmaps depicting two eigenfunctions $\phi_{\ell}$ of the kernel integral operator $\bm{K}_{N}$ from (8) for each of the three trajectories used as our training data. The top row shows the third eigenfunction for each of the three trajectories; the bottom row the fourth eigenfunction for each trajectory. From left to right, the first column corresponds to the trajectory data obtained for $\delta=0$, the second to $\delta=0.5$ and the third to $\delta=1$.
@@
Using the assumptions and methodology introduced in Section 4.5 for dealing with multitrajectory training data, we consider $\mu$ to be the measure generated by the full dynamics (12) restricted to the family of initial conditions (18) for $\delta\in[0,1]$. To sample this measure we use three trajectories generated using the values $\delta\in\{0,0.5,1\}$. For each trajectory we integrate the full dynamics (12) in time for 12,200 timesteps and then sample the next 3,260 timesteps. Each sample contains the values of functions $h$ and $q$ on the fine grid.
@@
To derive the resolved state samples that will be used as our training data, we coarsen the model and sampled data. By averaging the sampled data over every $M_{s}=20$ fine cells we arrive at a coarse grid of $M=96$ cells of length $\Delta\hat{s}=20\Delta s$ each. Given that each cell is now 202020 times longer we also increase the sampling timestep by the same amount to set $\Delta\hat{t}=20\Delta t$. From the original 3,260 samples per trajectory generated on the fine grid we keep every 20th sample, leading to 163 samples per trajectory. For each one of those samples we spatially average the solution over every 202020 cells to derive our resolved state samples, namely the averaged values of $h$ and $q$ over the coarse grid. We then derive the subgrid flux training samples using (17).
[…]
The number of delays $Q$ was chosen by a combination of physical reasoning and trial and error to balance the effectiveness of the derived basis in representing the true flux samples and the cost associated with its computation. A useful rule of thumb is to select a delay window that is comparable to the timescale of the physical processes of interest. In our case, we see that the traveling wave fronts interact approximately every 30 timesteps (1.5$\approx$ 1.5≈ 1.5 time units) for the solution in Figure 2 and approximately every 50 timesteps (2.5$\approx$ 2.5≈ 2.5 time units) for the solution in Figure 3, with similar results for additional solutions not shown here. Using $Q=64$ is sufficient to ensure that at least one period of this interaction is captured in each delay embedded sample. Additionally, increasing the value of $Q$ further did not lead to noticeable differences in our results. The stencil size $J$ used for the conditioning of quantum states was chosen empirically. Increasing the value of $J$ further was shown to slightly improve the performance of the bayesian conditioning operation, but with an associated increase in computational cost.
@@
To form the orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ we must compute the eigenvectors of an $NM\times NM$ symmetric kernel matrix $\bm{K}_{N}$ representing the data driven integral operator (8). To reduce the computational cost of this operation we perform a partial Cholesky factorization of $\bm{K}_{N}$ to approximate it by a lower rank matrix $\tilde{\bm{K}}_{N}=\bm{F}\bm{F}^{T}$ with $\bm{F}\in\mathbb{R}^{NM\times r}$ and $r<NM$. We perform the factorization using rank parameter $r=L=6144$ and the randomized algorithm developed in [66]. We then use the approximate kernel matrix to perform its bistochastic normalization and solve the eigenvalue problem with reduced cost [67].
A selection of the resulting eigenfunctions is presented in Figure 1, where we show two eigenfunctions for each of the three trajectories used to build our training dataset. The shown eigenfunctions exhibit propagating fronts which resemble the main features of the true state and flux samples presented in Figures 2 and 3. More specifically, they are approximately constant on the characteristic lines of the propagating wave fronts, as expected from their symmetry invariance property. Similar features are present in the rest of the eigenfunctions not shown here. As explained in Sections 4.2 and 4.3 the strong dynamic relevance of each eigenfunction is a result of its spatial symmetry invariance and of the fact that generally it cannot be represented as the tensor product of a single pair of spatial and temporal modes. These properties allow us to compute a compressed representation of the true dynamics with a moderate number of eigenfunctions, enhancing the computational efficiency of the resulting closure scheme.
@@
To test the predictive performance of the QMCl scheme we use two initial conditions obtained from (18) for $\delta\in\{0.25,0.75\}$ which are trajectories not included in the training dataset. After integrating the true dynamics (12) for 12,200 timesteps we spatially average the found solutions to generate the initial conditions that will be used to initialize the resolved state in QMCl. Using those initial conditions we integrate in time the coarse dynamics (15) using timestep $\Delta\hat{t}$ and the subgrid fluxes predicted by QMCl. In what follows we present results obtained after integrating for a total of 120 timesteps while performing bayesian conditioning of the quantum states every 10 timesteps.
Figure 2: Space-time heatmaps for the trajectory with $\delta=0.25$. The top row shows the results for the resolved state variables; the bottom row for the subgrid fluxes. For each row, from left to right: the first two plots show the results for the height field $h$, with the true results on the left and the predicted results on the right; the next two plots show the results for the momentum field $q$, with true results on the left and predicted on the right.
Figure 3: Space-time heatmaps for the trajectory with $\delta=0.75$. The layout of the plots is as in Figure 2.
@@
The results comparing the evolution under the true and surrogate dynamics are grouped into four figures. Figures 2 and 4 include the results for the trajectory generated by the initial condition obtained from (18) for $\delta=0.25$. Figures 3 and 5 showcase the results obtained for $\delta=0.75$.
Figure 4: Spatial profiles of the final solution snapshot for the trajectory with $\delta=0.25$. The top row shows the results for the resolved state and the subgrid fluxes for the height field $h$; the bottom row for the momentum field $q$. In each row, the true results are on the left subplot, the predicted ones on the right subplot. Within each subplot, full lines are used for the resolved state with their numerical values indicated on the left axis; dashed lines are used for the corresponding fluxes with their values indicated on the right axis.
Figure 5: Spatial profiles of the final solution snapshot for the trajectory with $\delta=0.75$. The layout of the plots is as in Figure 4.
Figure 6: Space-time heatmaps for the trajectory with $\delta=0.25$ comparing the dynamics on the coarse grid with the surrogate dynamics obtained by QMCl. From left to right, the first two plots show the results for the height field $h$, with the coarse grid results on the left and the QMCl results on the right; the next two plots show the results for the momentum field $q$, with coarse grid results on the left and QMCl on the right. The colorbars for each subplot are as in Figure 2.
@@
Looking at the heatmaps for the height and momentum fields in Figure 2 for the trajectory with $\delta=0.25$ and in Figure 3 for the one with $\delta=0.75$, we see that QMCl clearly captures the main features of the dynamics for the tested time horizon. For both trajectories QMCl correctly predicts the main traveling wave features of the dynamics, including the spatial variations of the waves as they interact with one another. Moreover, the magnitude of the predicted fields is consistent with the true dynamics in most regions of the spatial domain. In addition to the results shown in these figures, we have also verified that QMCl can reproduce the dynamical trajectories included in its training dataset with good accuracy.
@@
In the present closure model the subgrid fluxes play the role of antidiffusive corrections to the fluxes generated entirely on the coarse grid [80]. Their values are proportional to the spatial derivative of the associated field, taking large values when the height or momentum field is changing rapidly in space, and approximately zero value when the fields are constant in space. As a result, the contribution of the subgrid fluxes is critical in regions of large spatial variation of the height and momentum fields, counterbalancing the increased diffusion of the solution introduced by our coarsening of the original dynamics by spatial averaging. The strong spatial variation of the true subgrid fluxes renders the considered dynamics challenging to resolve and predict accurately by any closure scheme (Section 5.4). Typically, when the exact value of the subgrid fluxes is not predicted correctly by the employed closure model (derived by QMCl or another method), the surrogate evolution will be more diffusive than the true evolution.
@@
The effect of diffusion can be seen in the heatmaps for the two trajectories. By comparing the true and predicted fields of height and momentum, we see that the surrogate dynamics leads to stronger diffusion of the solution in areas where there is strong spatial variation. The increased diffusion is seen in the plots as bleeding of color between different space-time regions, with the boundaries between those regions being less sharp than in the true dynamics. Similarly, the same effect can be seen by comparing the subgrid flux fields for height and momentum. In the true dynamics, the peaks of the fluxes are sharp and clearly separated from the valleys, with clearly visible waves traveling in opposite directions. In the predicted dynamics, although the waves are still predicted correctly, their peaks are not as sharp, with their values diffusing more strongly to the surrounding region.
@@
The flux plots also clearly demonstrate the effect of bayesian conditioning, which for these results is performed every 101010 timesteps (0.5$\approx$ 0.5≈ 0.5 time units). One can see the discontinuities that occur every 101010 timesteps for the predicted fluxes over large regions of the spatial domain. These discontinuities represent corrections introduced to the predicted flux values every time conditioning of the quantum states is performed. We perform the bayesian conditioning every 101010 timesteps to reduce the computational cost of the simulation, since the conditioning is one of the most expensive operations in our closure scheme (Section 5.4). In general, the prediction accuracy increases with increased conditioning frequency. Based on our numerical experiments, conditioning the quantum states every 101010 timesteps strikes a good balance between prediction accuracy and computational cost for the considered dynamics.
@@
To further demonstrate the crucial role played by the subgrid fluxes in counterbalancing diffusion, in Figure 6 we compare the evolution of the QMCl surrogate dynamics with the true dynamics generated entirely on the coarse grid, namely with the subgrid fluxes (17) set to zero at every grid point. The comparison is carried out for the trajectory generated from the initial condition (18) with $\delta=0.25$ (Figures 2 and 4), but similar conclusions can be drawn from other initial conditions. The comparison demonstrates that the surrogate fluxes predicted by QMCl successfully counteract the diffusive effect of the spatial coarsening seen in the coarse grid simulation for the tested time horizon. Namely, although the subgrid fluxes predicted by QMCl are not as accurate as the true subgrid fluxes in regions of strong spatial variation of the solution, they are nevertheless effective in counterbalancing diffusion and thus improving the accuracy of the predicted dynamics.